OpenStudy (anonymous):

what is 4x/5+3/2 =2x??

OpenStudy (mathmagician):

4x+15/2=10x 6x=15/2 x=15/12

OpenStudy (anonymous):

thank you i was stuck in the middle with this

OpenStudy (anonymous):

Yeah, just move all your x terms to one side of the = and non-x terms to the other side, then divide by the coefficient in front of the x.

OpenStudy (anonymous):

thanks